# Shahriar Noroozizadeh

### v2

@@ 5.3 Failure of end-to-end composition learning
While $\ell$-fold composition functions are surprisingly easy to express in transformers [149], empirical results have time and again demonstrated that they are hard to learn through gradient-based methods, both in traditional deep network settings [156, 2, 61, 54, 1, 2] and more recently in language models too [125, 98, 34, 134, 194, 143, 34, 71, 160]. Others [13, 72, 157] demonstrate how next-token learning can trap training at a stage where composition learning becomes a problem. Theoretical works have attempted to formalize these failures by demonstrating limits due to expressivity [108, 32, 132] or sample complexity [156] or computational complexity [181, 72] or in terms of statistical queries [180].
@@
(Theoretical hardness of learning compositions) Existing theoretical hardness results typically only show that some worst-case function in a class of compositional functions is hard-to-learn; no proclamations are made about how hard it is to learn a fixed function we care about. Indeed, with contrived initial conditions, a singleton function class can be provably learned [4, 3, 118]. Yet, in practice, learning a fixed function is hard, proving which has been an open question—until the recent negative results of Abbe et al. [6], Shoshani and Shamir [159]. These show how the (fixed) full parity function is hard to learn with gradient descent which, through a reduction, proves that in-context composition tasks are hard to learn [72]. We leave it as an open question to link our in-weights composition task to one of these hardness results.
@@ 5.4 In-context graph tasks
While Bachmann and Nagarajan [13] identify the path-star topology as a failure case for next-token learning, Frydenlund [46, 47] demarcate the extent of this failure in the same in-context setting, whereas Brinkmann et al. [24] report positive path-finding results in other graph topologies. Others [150, 148, 190] study Transformers in various in-context graph tasks like search and counting. Connections between in-context graph tasks and spectral biases exist [35, 129] but should not be confused with the spectral bias in in-weights tasks. While all these works study symbolic graph tasks, other works have empirically identified the limitations on graphs described in natural language [63, 175, 37]. Kim et al. [86], Ying et al. [193] propose algorithmic ideas for encoding graphs as inputs to Transformers. Finally, various failures [114, 39, 168, 169, 170, 158] have been reported on in-context tasks, including planning tasks, framed as word problems.
@@
The dichotomy between drawing information from context vs. drawing information from the weights has been studied in various angles. Some have looked at this from the aspect of two competing circuits relying on one source vs. the other [29, 28, 121, 33]. Others have looked at it in the context of the learning paradigms of in-context learning and finetuning the weights [116, 90]. Closer to us, the stark in-context vs. in-weights disparity when it comes to handling global relationships has been emphasized in Wang et al. [174], Geerts et al. [48], for which our results provide further evidence.
@@ 5.5 Analyses of graph and word embedding methods
Much attention has been given to characterizing what embeddings are learned by various contrastive losses such as Node2Vec. Most of these are on losses simpler than the cross-entropy loss, with the exception of a recent line of work [199, 198, 167] which studies the geometry of cross-entropy-trained (weight-untied) Word2Vec models. However, certain assumptions are made, such as freezing one of the layers [167] or heuristically approximating via a regularized proxy model [199]. Curiously, [199] find that their system only directionally convergences, while the weights themselves diverge; we however point out that it’s reasonable to expect the model to converge to a zero-gradient solution. Regardless, [199] suggest an insightful connection to the converged direction to support vector machines; this may inspire better characterizations of what solution is found by our unregularized Node2Vec system.
The connection to a graph spectrum has been made in many other analyses. These analyses focus on a simpler loss called the negative sampling loss where the closed form expression for the inner products is straightforward (namely, the so-called pointwise mutual information (PMI) matrix, as discovered in Levy and Goldberg [93]). This analysis does not however tell us what exactly the embeddings are. The connection between these embeddings and the graph spectrum has been established in adjacent settings, like with DeepWalk [137] (where the objective is explicitly multi-hop), in low-rank settings like SimCLR [66, 165] or with quadratic losses with early stopping [83] or the softmax loss with rank $1$ [76]. Other analyses of Node2Vec focus on specific types of graphs such as stochastic block models [38, 67]. We refer the reader to Goyal and Ferrara [56] for a survey of graph embedding methods. Finally, we clarify that these methods and our insights must not be confused with graph neural networks [192, 200], where the graphs are not stored in parametric memory, but presented as input.
@@ 5.6 Other foundational works on generalization and memorization
Spectral and simplicity bias. The type of spectral bias we study in memorization and sequence modeling must be distinguished from the one studied in generalization and traditional classification and regression settings [140, 184, 82, 12, 146, 19]. In these earlier studies, the spectrum is that of a continuous function or a decision boundary (e.g., say, the Fourier components of a polynomial), whereas the spectrum we are concerned with is of a discrete, combinatorial object, namely the graph adjacency matrix. This corresponds to a difference in the nature of representations as remarked below.
@@
Memorization. Our work is also orthogonal to the seminal works of Zhang et al. [195], Neyshabur et al. [122] who were concerned with classical generalization tasks that possess statistical redundancies. Their argument is that explicit pressures cannot explain why representations arise in such tasks, but implicit pressures may. Our memorization task on the other hand is in sequence modeling, and lacks statistical redundancies; both explicit and implicit pressures do not suffice to explain the geometric representations. Our work is also orthogonal to the foundational work of Feldman [42] who argue that memorizing the quirks of a training set can be necessary for generalization in long tail datasets.
Linear representation hypothesis. A long-studied geometric concept in language models is the concept of linear representations in analogies [130, 131, 40, 112]. The introduction of certain concepts often takes a linear direction, surprisingly, independent of context i.e., going from cow to calf takes the same direction as cat to kitten, independent of the source in the context, (cow, cat). This linear structure, is related, but neither reducible to nor reducible from geometric memorization. Many theories have been proposed to model the linear geometry and semantics of these embeddings [53, 7, 41, 8, 68, 79]. These studies are orthogonal since their contribution lies in identifying what structures exist in word-context relationships for such geometries to arise in the embeddings. This is akin to identifying structures in the adjacency matrix, while our analysis is agnostic to such structures. There are many other complementary analyses of these embeddings, both theoretical [58] and empirical [113, 31, 30, 94].
@@ 7 Conclusions and Future Work
Various practical questions arise out of these findings. The elegant geometries of Node2Vec models indicate that Transformer memory can be made much more geometric. It is also unclear whether Transformers exhibit such desirable behaviors in more complex graph topologies. Importantly, empirical works on implicit reasoning in natural language have so far been mixed. More careful empirical research and ideation may be needed to make the geometric view more broadly applicable. Orthogonally, our findings are of relevance to making choices between parametric vs. contextual memory, and also between generative retrieval vs. dual encoder retrieval models.
@@
Acknowledgments. We would like to thank Gaurav Ghosal, Gintare Karolina Dziugaite, George H Chen, Christina Baek, Jacob Springer and Omar Salemohamed for valuable feedback on earlier versions of this draft. We are also grateful to Andrej Risteski, Alberto Bietti, Ekdeep Singh Lubana and Andrew Lampinen for discussions and pointers to related work.
@@ Appendix A Detailed background on the path-star task
Figures˜3 and 2 contrast the two evaluation regimes we use throughout: in-weights (edge memorization & path finetuning) versus in-context (graph in the prompt).
@@ A.1 Failure of next-token learning in the in-context path-star task
The Clever Hans Cheat. Early on during training, the model fits the later tokens in the target — concretely nodes $v_{2},\ldots,{v}_{\tt{goal}}$ — as the unique child of the previous token provided in the input during next-token training. This is known as a Clever Hans cheat: the model uses a local rule that relies on witnessing part of the ground truth prefix ($\bm{p},\bm{r}_{<i}$) to predict the next token $r_{i}$ — as against only using the prefix $\bm{p}$ to predict the answer tokens. Arguably, this happens because the cheat is much simpler to learn — as an induction head [126] — than the more complex solutions that rely only on $\bm{p}$ (which involves search-and-backtracking or the simpler lookahead-and-reverse). Simpler predictive features are prioritized early in training [12, 155, 147], which starves gradient signals from more complex features [133].
The Indecipherable Token. Therefore, in the second stage towards failure, the model attempts to learn the first token $r_{i}$—the key decision-making token—purely based on information about the graph in the prefix $\bm{p}$, and without any gradient signal about the later tokens. This is however a computationally hard “needle-in-the-haystack problem” [181]: the model needs to find a complex end-to-end algorithm with $\ell$ subroutines in it, without any intermediate supervision for those subroutines. The first token thus becomes “indecipherable” and the model simply memorizes it on the training data; during test-time the model subsequently predicts the first token as a random neighbor, and then continues following down the path by applying the local follow-the-child rule it learned through the Clever Hans cheat.
Fig.˜5 (also reproduced in Fig.˜13) demonstrates these failure modes empirically, showing that next-token prediction achieves only chance-level performance on path-star graphs of varying sizes, with the first token remaining particularly difficult to learn in isolation.
(Success of implicit in-weights reasoning) On in-weights path-star graphs of as many as $5\times 10^{4}$ nodes, trained on $75\%$ of the total $10^{4}$ paths, both the Transformer and Mamba are able to predict unseen paths when conditioned on held-out leaves with as much as 100% accuracy (see left plot of Fig.˜4 for Transformer (also reproduced in Fig.˜12), and left plot of Fig.˜14 for Mamba) Similar positive results for some harder graph topologies are in §C.2.
@@ A.1.1 Where does the path-star-failure argument go wrong in the in-weights setting?
Recall that the path-star failure unfolds in two stages. Perhaps one of these stages does not play out in the in-weights setting. A first possibility could be that the Clever Hans cheat is not picked up here. The cheat was a simple, left-to-right pattern that fits all but the first token as the unique neighbor of the preceding ground-truth token that was present as input. Such left-to-right cheats are simpler than the true right-to-left solution, and thus quickly learned. However, perhaps when the target we train on is an in-weights path, the cheat is not easy to learn—say, due to the nature of recalling from parametric memory. A complex cheat may not be learned quickly, allowing gradients from future tokens to reach the first token representation, in turn allowing the correct, right-to-left solution to compete and emerge. Indeed, such gradients, termed as pre-caching gradients [183] have been identified in a line of empirical works in natural language [127, 77, 145, 183, 110]. We summarize this hypothesis out below:
@@
Figure 12: Success of Transformer in in-weights path-star task. (left) A next-token-trained Transformer achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$ (Section˜A.1). (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success (Section˜A.1.1). (right) Success of hardest-token-only task. In fact, the hardest token (the first) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory view (§2.3). Analogous plots for Mamba are in Fig.˜14. We invite the reader to contrast these in-weights task results with the in-context task ones in Fig.˜13.
@@
We can indirectly test for this hypothesis as follows. If the model learned the right-to-left dependencies, it must also learn the tokens in the reverse order: the unique predecessor of the goal node is the easiest to identify (under the right-to-left dependencies) and will thus be learned first; the next-easiest is the next predecessor and so on, until the first node (which depends on all that has been learned so far). Indeed, in the in-context setting of B&N’24 where they explicitly switch off the Clever Hans cheat (under teacherless training), such a reverse-learning cascade is observed (Fig.˜5). However, we do not see this in the in-weights setting.
(No reverse-learning cascade) The target tokens in the in-weights path-star task are learned in no particular order by the Transformer and Mamba models (see middle plots of Figs.˜4 and 14 in contrast with Fig.˜5; or see Fig.˜35 for side-by-side comparison).
@@
Testing Section˜A.1.1 is easy: we train the model simply on the first token loss, instead of the loss over the full path sequence. We find that this task is trivial here, affirming Section˜A.1.1, and isolating a much stronger and cleaner instance of implicit in-weights reasoning.
@@ B.1 Graphs, tokens, and data construction
In-context datasets. Adapted from Bachmann and Nagarajan [13], the prefix contains a randomized adjacency serialization and $({v}_{\tt{root}},{v}_{\tt{goal}})$; the target is the full path. Where NTP fails in-context, we use the teacherless objective of Bachmann and Nagarajan [13] for comparison plots (details in §B.3).
@@ B.2.1 In-weights path-star task experiments
Backbone. The main experiments appearing in §2 use a decoder-only Transformer (GPT-mid) with a causal mask, pre-norm LayerNorm, sinusoidal positional embeddings [172], GELU MLPs, and tied input/output token embeddings. Additionally, experiments in §C.1 employ a Mamba sequence model [60] of comparable scale.

### v3

@@ 5.3 Failure of end-to-end composition learning
While $\ell$-fold composition functions are surprisingly easy to express in transformers [159], empirical results have time and again demonstrated that they are hard to learn through gradient-based methods, both in traditional deep network settings [166; 2; 67; 59; 1; 2] and more recently in language models too [135; 106; 38; 144; 206; 154; 38; 77; 170]. Others [14; 78; 167] demonstrate how next-token learning can trap training at a stage where composition learning becomes a problem. Theoretical works have attempted to formalize these failures by demonstrating limits due to expressivity [117; 35; 142] or sample complexity [166] or computational complexity [192; 78] or in terms of statistical queries [191].
@@
(Theoretical hardness of learning compositions) Existing theoretical hardness results typically only show that some worst-case function in a class of compositional functions is hard-to-learn; no proclamations are made about how hard it is to learn a fixed function we care about. Indeed, with contrived initial conditions, a singleton function class can be provably learned [4; 3; 127]. Yet, in practice, learning a fixed function is hard, proving which has been an open question—until the recent negative results of Abbe et al. [6]; Shoshani and Shamir [169]. These show how the (fixed) full parity function is hard to learn with gradient descent which, through a reduction, proves that in-context composition tasks are hard to learn [78]. We leave it as an open question to link our in-weights composition task to one of these hardness results.
@@ 5.4 In-context graph tasks
While Bachmann and Nagarajan [14] identify the path-star topology as a failure case for next-token learning, Frydenlund [50, 51] demarcate the extent of this failure in the same in-context setting, whereas Brinkmann et al. [26] report positive path-finding results in other graph topologies. Others [160; 158; 201] study Transformers in various in-context graph tasks like search and counting. Connections between in-context graph tasks and spectral biases exist [39; 139] but should not be confused with the spectral bias in in-weights tasks. While all these works study symbolic graph tasks, other works have empirically identified the limitations on graphs described in natural language [69; 185; 41]. Kim et al. [92]; Ying et al. [204] propose algorithmic ideas for encoding graphs as inputs to Transformers. Finally, various failures [123; 43; 178; 179; 180; 168] have been reported on in-context tasks, including planning tasks, framed as word problems.
@@
The dichotomy between drawing information from context vs. drawing information from the weights has been studied in various angles. Some have looked at this from the aspect of two competing circuits relying on one source vs. the other [32; 31; 130; 36]. Others have looked at it in the context of the learning paradigms of in-context learning and finetuning the weights [125; 97]. Closer to us, the stark in-context vs. in-weights disparity when it comes to handling global relationships has been emphasized in Wang et al. [184]; Geerts et al. [53], for which our results provide further evidence.
@@ 5.5 Analyses of graph and word embedding methods
Much attention has been given to characterizing what embeddings are learned by various contrastive losses such as Node2Vec. Most of these are on losses simpler than the cross-entropy loss, with the exception of a recent line of work [211; 210; 177] which studies the geometry of cross-entropy-trained (weight-untied) Word2Vec models. However, certain assumptions are made, such as freezing one of the layers [177] or heuristically approximating via a regularized proxy model [211]. Curiously, [211] find that their system only directionally converges, while the weights themselves diverge; we however point out that it’s reasonable to expect the model to converge to a zero-gradient solution. Regardless, [211] suggest an insightful connection to the converged direction to support vector machines; this may inspire better characterizations of what solution is found by our unregularized Node2Vec system.
The connection to a graph spectrum has been made in many other analyses. These analyses focus on a simpler loss called the negative sampling loss where the closed form expression for the inner products is straightforward (namely, the so-called pointwise mutual information (PMI) matrix, as discovered in Levy and Goldberg [100]). This analysis does not however tell us what exactly the embeddings are. The connection between these embeddings and the graph spectrum has been established in adjacent settings, like with DeepWalk [148] (where the objective is explicitly multi-hop), in low-rank settings like SimCLR [72; 175] or with quadratic losses with early stopping [89] or the softmax loss with rank $1$ [82]. Other analyses of Node2Vec focus on specific types of graphs such as stochastic block models [42; 73]. We refer the reader to Goyal and Ferrara [61] for a survey of graph embedding methods. Finally, we clarify that these methods and our insights must not be confused with graph neural networks [203; 213], where the graphs are not stored in parametric memory, but presented as input.
@@ 5.6 Other foundational works on generalization and memorization
Spectral and simplicity bias. The type of spectral bias we study in memorization and sequence modeling must be distinguished from the one studied in generalization and traditional classification and regression settings [151; 195; 88; 13; 19; 21]. In these earlier studies, the spectrum is that of a continuous function or a decision boundary (e.g., say, the Fourier components of a polynomial), whereas the spectrum we are concerned with is of a discrete, combinatorial object, namely the graph adjacency matrix. This corresponds to a difference in the nature of representations as remarked below.
@@
Memorization. Our work is also orthogonal to the seminal works of Zhang et al. [207]; Neyshabur et al. [131] who were concerned with classical generalization tasks that possess statistical redundancies. Their argument is that explicit pressures cannot explain why representations arise in such tasks, but implicit pressures may. Our memorization task on the other hand is in sequence modeling, and lacks statistical redundancies; both explicit and implicit pressures do not suffice to explain the geometric representations. Our work is also orthogonal to the foundational work of Feldman [46] who argue that memorizing the quirks of a training set can be necessary for generalization in long tail datasets.
Linear representation hypothesis. A long-studied geometric concept in language models is the concept of linear representations in analogies [140; 141; 44; 121]. The introduction of certain concepts often takes a linear direction, surprisingly, independent of context i.e., going from cow to calf takes the same direction as cat to kitten, independent of the source in the context, (cow, cat). This linear structure, is related, but neither reducible to nor reducible from geometric memorization. Many theories have been proposed to model the linear geometry and semantics of these embeddings [58; 7; 45; 8; 74; 85]. These studies are orthogonal since their contribution lies in identifying what structures exist in word-context relationships for such geometries to arise in the embeddings. This is akin to identifying structures in the adjacency matrix, while our analysis is agnostic to such structures. There are many other complementary analyses of these embeddings, both theoretical [63] and empirical [122; 34; 33; 101].
@@ 7 Conclusions and Future Work
Various practical questions arise from these findings. The elegant geometries of Node2Vec models indicate that Transformer memory can be made much more geometric. It is also unclear whether Transformers exhibit such desirable behaviors in more complex graph topologies. Importantly, empirical works on implicit reasoning in natural language have so far been mixed. More careful empirical research and ideation may be needed to make the geometric view more broadly applicable. Orthogonally, our findings are of relevance to making choices between parametric vs. contextual memory, and also between generative retrieval vs. dual encoder retrieval models.
@@
Acknowledgments. We would like to thank Gaurav Ghosal, Gintare Karolina Dziugaite, George H Chen, Christina Baek, Jacob Springer for valuable feedback on earlier versions of this manuscript. We are also deeply grateful to Omar Salemohamed, Andrej Risteski, Alberto Bietti, Ekdeep Singh Lubana and Andrew Lampinen for many discussions and pointers to related work.
@@ Appendix A Detailed background on the path-star task
Figures˜10 and 11 contrast the two evaluation regimes we use throughout: in-weights (edge memorization & path finetuning) versus in-context (graph in the prompt).
@@ A.1 Failure of next-token learning in the in-context path-star task
The Clever Hans Cheat. Early on during training, the model fits the later tokens in the target — concretely nodes $v_{2},\ldots,{v}_{\tt{goal}}$ — as the unique child of the previous token provided in the input during next-token training. This is known as a Clever Hans cheat: the model uses a local rule that relies on witnessing part of the ground truth prefix ($\bm{p},\bm{r}_{<i}$) to predict the next token $r_{i}$ — as against only using the prefix $\bm{p}$ to predict the answer tokens. Arguably, this happens because the cheat is much simpler to learn — as an induction head [Olsson et al., 2022] — than the more complex solutions that rely only on $\bm{p}$ (which involves search-and-backtracking or the simpler lookahead-and-reverse). Simpler predictive features are prioritized early in training [Arpit et al., 2017, Shah et al., 2020, Rosenfeld and Risteski, 2024], which starves gradient signals from more complex features [Pezeshki et al., 2021].
The Indecipherable Token. Therefore, in the second stage towards failure, the model attempts to learn the first token $r_{i}$—the key decision-making token—purely based on information about the graph in the prefix $\bm{p}$, and without any gradient signal about the later tokens. This is however a computationally hard “needle-in-the-haystack problem” [Wies et al., 2023]: the model needs to find a complex end-to-end algorithm with $\ell$ subroutines in it, without any intermediate supervision for those subroutines. The first token thus becomes “indecipherable” and the model simply memorizes it on the training data; during test-time the model subsequently predicts the first token as a random neighbor, and then continues following down the path by applying the local follow-the-child rule it learned through the Clever Hans cheat.
Fig.˜13 demonstrates these failure modes empirically, showing that next-token prediction achieves only chance-level performance on path-star graphs of varying sizes, with the first token remaining particularly difficult to learn in isolation.
(Success of implicit in-weights reasoning) On in-weights path-star graphs of as many as $5\times 10^{4}$ nodes, trained on $75\%$ of the total $10^{4}$ paths, both the Transformer and Mamba are able to predict unseen paths when conditioned on held-out leaves with as much as 100% accuracy (see left plot of Fig.˜12 for Transformer), and left plot of Fig.˜14 for Mamba). Similar positive results for some harder graph topologies are in §C.2.
@@ A.1.1 Where does the path-star-failure argument go wrong in the in-weights setting?
Recall that the path-star failure unfolds in two stages. Perhaps one of these stages does not play out in the in-weights setting. A first possibility could be that the Clever Hans cheat is not picked up here. The cheat was a simple, left-to-right pattern that fits all but the first token as the unique neighbor of the preceding ground-truth token that was present as input. Such left-to-right cheats are simpler than the true right-to-left solution, and are thus quickly learned. However, perhaps when the target we train on is an in-weights path, the cheat is not easy to learn—say, due to the nature of recalling from parametric memory. A complex cheat may not be learned quickly, allowing gradients from future tokens to reach the first token representation, in turn allowing the correct, right-to-left solution to compete and emerge. Indeed, such gradients, termed as pre-caching gradients [Wu et al., 2024] have been identified in a line of empirical works in natural language [Pal et al., 2023, Jenner et al., 2024, Rofin et al., 2025, Wu et al., 2024, Men et al., 2024]. We summarize this hypothesis out below:
@@
Figure 12: Success of Transformer in in-weights path-star task. (§2) (left) A next-token-trained Transformer achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$. (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success. (right) Success of hardest-token-only task. In fact, the hardest token (the first) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory view (§2.3). Analogous plots for Mamba are in Fig.˜14. We invite the reader to contrast these in-weights task results with the in-context task ones in Fig.˜13.
@@
We can indirectly test for this hypothesis as follows. If the model learned the right-to-left dependencies, it must also learn the tokens in the reverse order: the unique predecessor of the goal node is the easiest to identify (under the right-to-left dependencies) and will thus be learned first; the next-easiest is the next predecessor and so on, until the first node (which depends on all that has been learned so far). Indeed, in the in-context setting of B&N’24 where they explicitly switch off the Clever Hans cheat (under teacherless training), such a reverse-learning cascade is observed (Fig.˜13). However, we do not see this in the in-weights setting.
(No reverse-learning cascade) The target tokens in the in-weights path-star task are learned in no particular order by the Transformer and Mamba models (see middle plots of Figs.˜12 and 14 in contrast with Fig.˜13).
@@
Testing Section˜A.1.1 is easy: we train the model simply on the first token loss, instead of the loss over the full path sequence. We find (as discussed in the main paper) that this task is successful here, affirming Section˜A.1.1, and isolating a much stronger and cleaner instance of implicit in-weights reasoning.
@@ B.1 Graphs, tokens, and data construction
In-context datasets. Adapted from B&N’24, the prefix contains a randomized adjacency serialization and $({v}_{\tt{root}},{v}_{\tt{goal}})$; the target is the full path. Where NTP fails in-context, we use the teacherless objective of B&N’24 for comparison plots (details in §B.3).
@@ B.2.1 In-weights path-star task experiments
Backbone. The main experiments appearing in §2 use a decoder-only Transformer (GPT-mid) with a causal mask, pre-norm LayerNorm, sinusoidal positional embeddings [Vaswani et al., 2017], GELU MLPs, and tied input/output token embeddings. Additionally, experiments in §C.1 employ a Mamba sequence model [Gu and Dao, 2023] of comparable scale.